# Causally constrained reduced-order neural models of complex turbulent dynamical systems

### v2

@@ VIII Response to large step function forcings in the Charney-DeVore model
We now consider the nonlinear response regime by adding a step function forcing $\mathbf{F}$ to both the numerical (CdV) and neural models. In the main text we focused on the case $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std. dev. of the $x_{1}$ variable. We stress that this is a very large forcing, imposed on a model that had only access to unperturbed variability. Here, we perform six forcing experiments, applying a step forcing to each degree of freedom in turn with amplitude equal to its standard deviation: $\mathbf{F}=(\sigma_{1},0,\ldots,0)$, $(0,\sigma_{2},0,\ldots,0)$, $\ldots$, $(0,\ldots,0,\sigma_{6})$. For each case, we analyze the time-dependent response of the ensemble mean and variance of the full system. Results are shown in Fig. 7 (mean) and Fig. 8 (variance). The causally coarse-grained emulator outperforms the unconstrained, vanilla emulator. Notably, the vanilla emulator becomes unstable in three out of six forcing experiments (rows 1, 2, and 6 in Fig. 7), whereas the causally constrained emulator remains stable in all cases. Errors in the ensemble-mean response are primarily observed when the step forcing is applied to $x_{5}$ and $x_{6}$. To elucidate the origin of these errors, we revisit the response operator $R_{k,j}(t)$ inferred in the previous sections and focus on the ground-truth response estimated from the numerical model. We quantify the cumulative sensitivity of the system to perturbations applied to each degree of freedom $x_{j}$ as $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$,; where is $\tau_{\infty}=40$ Model Time Units, consistent with all response analyses. The resulting sensitivities are shown in Fig. 9. This analysis identifies $x_{5}$ and $x_{6}$ as the most sensitive degrees of freedom: perturbations applied to these variables induce the largest cumulative system-wide response. This high sensitivity provides a plausible, qualitative explanation for the observed degradation in ensemble-mean responses under large forcings applied to $x_{5}$ and $x_{6}$, where even small modeling errors are strongly amplified.
@@
Figure 7: Perturbed statistics in the nonlinear regime. Row (i) shows the repsonse in ensemble mean to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean response a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
@@
Figure 8: Perturbed statistics in the nonlinear regime. Row (i) shows the response in ensemble variance to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean to a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Responses are computed using an ensemble of $N_{e}=10^{5}$ members. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
Figure 9: Cumulative (causal) sensitivity $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$ of the Charney-DeVore model in Eq. (3) in the main text. For the response operator $R_{k,j}(t)$ we use the correct ground-truth. The system is more sensitive to perturbations on degrees of freedom $x_{5}$ and $x_{6}$, therefore providing a possible explanation for the degradation of results when adding large, step function forcings on these degrees of freedoms.

### v3

@@ IX Response to large step function forcings in the Charney-DeVore model
We now consider the nonlinear response regime by adding a step function forcing $\mathbf{F}$ to both the numerical (CdV) and neural models. In the main text we focused on the case $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std. dev. of the $x_{1}$ variable. We stress that this is a very large forcing, imposed on a model that had only access to unperturbed variability. Here, we perform six forcing experiments, applying a step forcing to each degree of freedom in turn with amplitude equal to its standard deviation: $\mathbf{F}=(\sigma_{1},0,\ldots,0)$, $(0,\sigma_{2},0,\ldots,0)$, $\ldots$, $(0,\ldots,0,\sigma_{6})$. For each case, we analyze the time-dependent response of the ensemble mean and variance of the full system. Results are shown in Fig. 10 (mean) and Fig. 11 (variance). The emulator constrained by both physics and causality generally leads to a better representation of the responses. This is clearly shown in the total MSE in Figure 12, aggregated over all response and perturbations (12), where the “physics and causal” emulator shows lower error in the long time response.
@@
Figure 10: Perturbed statistics in the nonlinear regime. Row (i) shows the repsonse in ensemble mean to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean response a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
@@
Figure 11: Perturbed statistics in the nonlinear regime. Row (i) shows the response in ensemble variance to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean to a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Responses are computed using an ensemble of $N_{e}=10^{5}$ members. Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
Figure 12: Perturbed statistics in the nonlinear regime. Panel (a): Total MSE over time for the ensemble mean response. Panel (b): Same as Panel (a) but the response in ensemble variance.
X Symmetry-Broken Lorenz-96 (L96) system
Random forcing $F_{j}$.
In this work, we introduce a time-independent forcing in the L96 system defined as $F_{j}=F+\eta_{j}$, where $\eta_{j}\sim\mathcal{N}(0,\sigma_{F}^{2})$ with $\sigma_{F}=10$. Due to the random component, we report the specific realization used in our experiments for reproducibility: $F_{j}=$ (6.1087865 , 12.32213349, 28.87925261, 17.93974419, 25.202309 , 21.77103791, 9.63536354, 21.4195222 , 12.83404549, 12.77610884, 16.97167319, 4. , 27.92166104, 9.28910325, 26.0026942 , 17.36321124, 31.3203308 , 9.40030586, 12.88205144, 19.37769127).
X.1 Stationary statistics
The marginals of the invariant distribution of the L96 system predicted by the two emulators are shown in Fig. 13. The corresponding autocorrelation functions are reported in Fig. 14.
Figure 13: Stationary statistics. Stationary distributions (PDFs) of the variables $x_{1},...,x_{20}$ of the L96 system comsidered in the main text. Labels: “L96” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
Figure 14: Autocorrelation functions (ACFs) of the variables $x_{1},...,x_{20}$ of the L96 system comsidered in the main text. Labels: “L96” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.